OpenStudy (anonymous):

Your friend hands you a graph of the engine power statistics of a race car. He says, “I know this graph is f(x) = –3(x + 4)6 – 8 but I can’t remember how it is related to the graph of x6.” Explain to your friend how the graph f(x) is a translation of the graph x6.

OpenStudy (anonymous):

can u help with mines first I will medal u

OpenStudy (anonymous):

I really need help first, can you help me

OpenStudy (anonymous):

@saynabahmed13

OpenStudy (anonymous):

I am a 6th grader is this 6th grade math

OpenStudy (anonymous):

Just to make sure what your equations are, \[f(x) = -3(x+4)^6-8\] and \[f(x) = x^6\] hint:Reflection